# Asymptotic Stability of 3D Out-flowing Compressible Viscous Fluid under Non-Spherical Perturbation

### v1

@@ 3.2 Estiamtes for Temporal Derivatives
First, we remark that the temporal derivative preserves the boundary condition (1.2), in the sense that
$\partial_{t}\psi(t,x)=\partial_{t}^{2}\psi(t,x)=0\qquad\text{for }\ (t,x)\in\Gamma_{T}\equiv[0,T]\times\partial\Omega.$ (3.24)
@@
There exists a positive constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for all $T\in[0,T_{0}]$ and $k=0,1$,
@@
Taking $\partial_{t}$ on (3.8a), then multiplying the resultant equality with $\frac{q_{+}}{\rho_{+}}\partial_{t}\phi$, we get
$\partial_{t}\big{(}\frac{q_{+}}{2\rho_{+}}|\partial_{t}\phi|^{2}\big{)}+\textnormal{{div}}\big{(}\frac{|\partial_{t}\phi|^{2}}{2}u\big{)}+q_{+}\partial_{t}\phi\textnormal{{div}}\partial_{t}\psi$
@@
In addition, taking $\partial_{t}$ on (3.8b), multiplying the resultant equality with $\partial_{t}\psi$, and integrating by parts, we get
@@
Adding (3.25)–(3.26), integrating the resultant equality in $(t,x)\in Q_{T}$, applying the boundary condition (3.24), Lemma 1.1, and inequality (3.12), we obtain
@@
Next, taking $\partial_{t}^{k}$ on (3.8a), then multiplying the resultant equality with $\partial_{t}^{k+1}\phi$, we get
@@
Set $k=0,1$. Applying $\partial_{t}^{k}$ on (3.8b), then taking inner-product on the resultant equality with $\partial_{t}^{k+1}\psi$, we get
@@
Adding (3.28)–(3.29), integrating in $(t,x)\in Q_{T}$, using boundary condition (3.24) and inequality (3.12), we get
@@
$=$ $\frac{\rho_{+}q_{+}^{2}}{2(\mu+\lambda)}\|\partial_{t}^{k}\phi(t,\cdot)\|^{2}\big{|}_{t=0}^{t=T}+\iint_{Q_{T}}\partial_{t}^{k+1}\psi\cdot\partial_{t}^{k}f$
@@
which holds true for $k=0,1$. Applying Lemma 3.4, inequality (3.12), and the previous estimate (3.27) on the above, we get
@@ 3.3 Estimates for Angular Derivatives
Consider the spherical coordinate system $(r,\theta,\varphi)$, where $r\geq 1$ is the radial variable, $\theta\in[0,\pi]$ is the polar angle, and $\varphi\in[0,2\pi)$ is the azimuthal angle. For the angular derivatives, terms such as $1/\sin\theta$ occur, which form singularity at $\theta=0$ or $\pi$. To circumvent this issue, we consider the following partition of unity: let
$\Omega_{V}\vcentcolon=\Big{\{}x=(x_{1},x_{2},x_{3})^{\top}\in\mathbb{R}^{3}\ \Big{|}\ |x|\geq 1,\quad\arccos\big{(}\frac{x_{3}}{|x|}\big{)}\in\big{[}\frac{\pi}{9},\frac{8\pi}{9}\big{]}\Big{\}},$
$\Omega_{H}\vcentcolon=\Big{\{}x=(x_{1},x_{2},x_{3})^{\top}\in\mathbb{R}^{3}\ \Big{|}\ |x|\geq 1,\quad\arccos\big{(}\frac{x_{2}}{|x|}\big{)}\in\big{[}\frac{\pi}{9},\frac{8\pi}{9}\big{]}\Big{\}}.$
Then $\Omega=\Omega_{V}\cup\Omega_{H}$. On each domain $\Omega_{V}$ or $\Omega_{H}$, we define the spherical coordinate variables $(r,\theta_{V},\varphi_{V})$ and $(r,\theta_{H},\varphi_{H})$ as follows:
$x_{1}=r\cos\varphi_{V}\sin\theta_{V},\quad$ $x_{2}=r\sin\varphi_{V}\sin\theta_{V},\quad$ $x_{3}=r\cos\theta_{V},\quad$ for $\text{for }x\in\Omega_{V},$
$x_{1}=r\cos\varphi_{H}\sin\theta_{H},\quad$ $x_{2}=r\cos\theta_{H},\quad$ $x_{3}=r\sin\varphi_{H}\sin\theta_{H},\quad$ for $\text{for }x\in\Omega_{H},$
@@
$\hat{r}\vcentcolon=\frac{x}{|x|},\quad\hat{\theta}_{V}\vcentcolon=\frac{x_{1}x_{3}\hat{e}_{1}+x_{2}x_{3}\hat{e}_{2}-(x_{1}^{2}+x_{2}^{2})\hat{e}_{3}}{|x|\sqrt{x_{1}^{2}+x_{2}^{2}}},\quad\hat{\varphi}_{V}\vcentcolon=\frac{-x_{2}\hat{e}_{1}+x_{1}\hat{e}_{2}}{\sqrt{x_{1}^{2}+x_{2}^{2}}}.$
For the case $x\in\Omega_{H}$, unit vectors $(\hat{r},\hat{\theta}_{H},\hat{\varphi}_{H})$ is defined by exchanging the terms $x_{2}\leftrightarrow x_{3}$ and $\hat{e}_{2}\leftrightarrow\hat{e}_{3}$ on the above expression. Let $\chi_{V}$, $\chi_{H}\in\mathcal{C}^{\infty}(\mathbb{R}^{3})$ be non-negative functions such that for all $x\in\mathbb{R}^{3}$, and $k=1,2,3,4$,
$0\leq\chi_{V},\,\chi_{H}\leq 1,\quad\chi_{V}+\chi_{V}\geq 1,\quad\text{supp}\chi_{V}\subseteq\Omega_{V},\quad\text{supp}\chi_{H}\subseteq\Omega_{H},$ (3.30a)
$|\nabla\chi_{H}|^{2}\leq C\chi_{H},\qquad|\nabla\chi_{V}|^{2}\leq C\chi_{V},\qquad|D^{k}\chi_{H}|+|D^{k}\chi_{V}|\leq C,$ (3.30b)
$\hat{r}\cdot\nabla\chi_{H}=\hat{\varphi}_{H}\cdot\nabla\chi_{H}=0,\qquad\hat{r}\cdot\nabla\chi_{V}=\hat{\varphi}_{V}\cdot\nabla\chi_{V}=0,$ (3.30c)
where $C>0$ is some generic constant. The detailed construction for $\chi_{V}$ and $\chi_{H}$ described above, can be found in Appendix B. Next, we give the precise definition for radial and angular derivatives. Consider the case $x\in\Omega_{V}$. We use the following conventions to denote the differential operators for spherical coordinate directions:
$\partial\in\{\partial_{\theta_{V}},\partial_{\varphi_{V}},\partial_{\theta_{H}},\partial_{\varphi_{H}}\},\qquad D\in\{\partial_{r},\partial_{\theta_{V}},\partial_{\varphi_{V}},\partial_{\theta_{H}},\partial_{\varphi_{H}}\},$
where
$\partial_{r}\vcentcolon=\frac{x}{|x|}\cdot\nabla,\qquad\partial_{\theta_{V}}\vcentcolon=|x|\hat{\theta}_{V}\cdot\nabla,\qquad\partial_{\varphi_{V}}\vcentcolon=\sqrt{x_{1}^{2}+x_{2}^{2}}\,\,\hat{\varphi}_{V}\cdot\nabla,$
$\partial_{\theta_{H}}\vcentcolon=|x|\hat{\theta}_{H}\cdot\nabla,\qquad\partial_{\varphi_{H}}\vcentcolon=\sqrt{x_{1}^{2}+x_{3}^{2}}\,\,\hat{\varphi}_{H}\cdot\nabla$
@@
$\chi\partial^{k}\psi|_{x\in\partial\Omega}=0\quad\text{for }\ \partial\in\{\partial_{\theta_{V}},\partial_{\varphi_{V}},\partial_{\theta_{H}},\partial_{\varphi_{H}}\}\ \text{ and }\ k=1,2,3.$ (3.31)
In what follows, we only derive estimates for one of $\Omega_{V}$ and $\Omega_{H}$, since the other case follows from the same argument. To simplify, we use the following notations,
$\chi\equiv\chi_{V}\ \text{or}\ \chi_{H},\qquad(\theta,\varphi)\equiv(\theta_{V},\varphi_{V})\ \text{or}\ (\theta_{H},\varphi_{H}),\qquad\tilde{\Omega}\equiv\Omega_{V}\ \text{or}\ \Omega_{H},$
$\tilde{Q}_{T}\equiv[0,T]\times\Omega_{V}\ \text{ or }\ [0,T]\times\Omega_{H},\qquad\tilde{\Gamma}_{T}\equiv[0,T]\times\partial\Omega_{V}\ \text{ or }\ [0,T]\times\partial\Omega_{H}.$
@@
There exists a constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for $\delta\in(0,1]$, $T\in[0,T_{0}]$, and $k=1,2,3$,
@@
Taking $\partial^{k}$ on equation (3.8a), and multiplying both sides of the resultant equality with $\chi\partial^{k}\phi$, we get
@@
Taking $\partial^{k}$ on equation (3.8b), and multiplying both sides of the resultant equality with $\chi\partial^{k}\psi$, we get
@@
Combining the two equations, integrating in $(t,x)\in Q_{T}\equiv[0,T]\times\Omega$, and integrating by parts using boundary condtion (3.31), we have
@@
By the property of cut-off function $\chi$ in (3.30), we have for all $\epsilon\in(0,1]$,
@@
Using the commutation relations, Proposition B.1, we also get for all $\epsilon\in(0,1]$,
@@
For the term $\mathcal{I}_{3}$, we recall the definition (3.8d). Then integrating by parts, applying Lemma 1.1 and inequality (3.12), we get for $k=1,2,3$,
@@
Applying Propositions A.1 and B.1 we obtain that
@@
Substituting $\mathcal{I}_{1}$$\mathcal{I}_{4}$ into (3.32), it holds that for $k=1,2,3$, and $\epsilon\in(0,1]$,
@@
$\leq$ $C\big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\big{\}}+\epsilon N^{2}(T)+\frac{C}{\epsilon}\sum_{m=0}^{k-1}\iint_{\tilde{Q}_{T}}|\nabla\partial^{m}\psi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t.$ (3.33)
Take $k=1$ and $\epsilon=\delta^{4}$ for some $\delta\in(0,1)$ in the above inequality. Then by Lemma 3.4,
@@
Next, take $k=2$ and $\epsilon=\delta^{2}$ in (3.33). Then applying Lemma 3.4, and substituting (3.34) into the resultant inequality, we get
@@
Finally, take $k=3$ and $\epsilon=\delta$ in (3.33). Then applying Lemma 3.4, and substituting (3.34)–(3.35) into the resultant inequality, we get
@@
There exists a constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for all $\delta\in(0,1]$, $T\in[0,T_{0}]$, and $k=0,1,2,3$,
@@
$\iint_{Q_{T}}\chi|\partial^{k}D_{t}\phi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t\leq\frac{C}{\delta^{7}}\Big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\Big{\}}+C\delta N^{2}(T).$
@@
The equation (1.1a) yields $D_{t}\phi=-\rho_{+}\textnormal{{div}}\psi+f^{0}$. Integrating this equation in $(t,x)\in Q_{T}$, applying Lemmas 3.4 and inequality (3.12), we obtain the first estimate. Next, taking $\partial^{k}$ on (3.8a) yields $\partial^{k}D_{t}\phi=-\rho_{+}\textnormal{{div}}\partial^{k}\psi-\rho_{+}[\partial^{k},\textnormal{{div}}]\psi+\partial^{k}f^{0}$. Integrating this equation in $(t,x)\in Q_{T}$, applying Lemmas 3.4, 3.6, Proposition B.1, and inequality (3.12), we obtain the second estimate. $\square$

### v2

@@ 2.3 Estiamtes for temporal derivatives
In this section, we obtain estimates for the temporal derivatives of $(\phi,\psi)$. We remark that the temporal derivative preserves the boundary condition (2.1a) in the sense that
$\partial_{t}\psi(t,x)=\partial_{t}^{2}\psi(t,x)=0\qquad\text{for }\ (t,x)\in\Gamma_{T}\vcentcolon=[0,T]\times\partial\Omega.$ (2.17)
@@
Suppose $(\phi,\psi)\in X(T_{0})$ is a solution to the system (2.1), where $T_{0}>0$ is the time given in (2.3)–(2.4). Then there exists a constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for $T\in[0,T_{0}]$ and $k=0,1$,
@@
Taking the derivative $\partial_{t}$ on (2.1b) and multiplying the resultant equality by $\frac{q_{+}}{\rho_{+}}\partial_{t}\phi$, we get
$\partial_{t}\big(\frac{q_{+}}{2\rho_{+}}|\partial_{t}\phi|^{2}\big)+\textnormal{{div}}\big(\frac{q_{+}}{2\rho_{+}}u|\partial_{t}\phi|^{2}\big)+q_{+}\partial_{t}\phi\textnormal{{div}}\partial_{t}\psi$
@@
In addition, taking the derivative $\partial_{t}$ on (2.1c), multiplying the resultant equality by $\partial_{t}\psi$, and integrating by parts, we get
@@
Summing equations (2.18)–(2.19), integrating the resultant equality in $(t,x)\in Q_{T}$, applying the boundary condition (2.17), Lemma 1.1, and inequality (2.5), we obtain
@@
Taking the derivative $\partial_{t}^{k}$ on (2.1b) for $k=0,1,$ and multiplying the resultant equality by $\partial_{t}^{k+1}\phi$, we obtain
@@
Taking the derivative $\partial_{t}^{k}$ on (2.1c) for $k=0,1$, and taking inner-product on the resultant equality with $\partial_{t}^{k+1}\psi$, we get
@@
Summing equations (2.21)–(2.22), integrating in $(t,x)\in Q_{T}$, using boundary condition (2.17) and inequality (2.5), we get
@@
$=$ $\frac{\rho_{+}q_{+}^{2}}{2(\mu+\lambda)}\|\partial_{t}^{k}\phi(t,\cdot)\|^{2}\big|_{t=0}^{t=T}+\iint_{Q_{T}}\partial_{t}^{k+1}\psi\cdot\partial_{t}^{k}f\,{\mathrm{d}}x{\mathrm{d}}t$
@@
which holds true for $k=0,1$. Applying Lemma 2.4, inequality (2.5) and the previous estimate (2.20) on the above equality, we get
@@ 2.4 Estimates for angular derivatives
Our aim is to obtain the estimates for spatial derivatives $\|(\nabla^{k}\psi,\nabla^{k}\phi)(t)\|$ where $k=1,2,3$. However the boundary condition (1.2) poses problem for the derivation of parabolic estimates using equations (2.1c). To resolve this issue, we decompose the gradient operator $\nabla$ into the tangential and normal differential operators, relative to the boundary surface $\partial\Omega$. Since the tangential differential operators preserves the boundary condition, it is viable to derive the parabolic estimates for the tangential derivatives of $(\psi,\phi)$. Thus the main purpose of this subsection is to specify the notion of tangential differential operators $\partial$ and obtain their corresponding estimates.
We consider the spherical coordinate system $(r,\theta,\varphi)$, where $r\geq 1$ is the radial variable, $\theta\in[0,\pi]$ is the polar angle and $\varphi\in[0,2\pi)$ is the azimuthal angle. For the angular derivatives, terms such as $1/\sin\theta$ occur, which form singularity at $\theta=0$ or $\pi$. To circumvent this issue, we employ the following partition of unity
$\Omega_{V}\vcentcolon=\Big\{x=(x_{1},x_{2},x_{3})^{\top}\in\mathbb{R}^{3}\ \Big|\ |x|\geq 1,\qquad\arccos\big(\frac{x_{3}}{|x|}\big)\in\big[\frac{\pi}{9},\frac{8\pi}{9}\big]\Big\},$
$\Omega_{H}\vcentcolon=\Big\{x=(x_{1},x_{2},x_{3})^{\top}\in\mathbb{R}^{3}\ \Big|\ |x|\geq 1,\qquad\arccos\big(\frac{x_{2}}{|x|}\big)\in\big[\frac{\pi}{9},\frac{8\pi}{9}\big]\Big\}.$
It is obvious that $\Omega=\Omega_{V}\cup\Omega_{H}$. On each domain $\Omega_{V}$ or $\Omega_{H}$, we define the spherical coordinate variables $(r,\theta_{V},\varphi_{V})$ and $(r,\theta_{H},\varphi_{H})$ as
$x_{1}=r\cos\varphi_{V}\sin\theta_{V},\qquad$ $x_{2}=r\sin\varphi_{V}\sin\theta_{V},\qquad$ $x_{3}=r\cos\theta_{V},\qquad$ $\text{for }x\in\Omega_{V},$
$x_{1}=r\cos\varphi_{H}\sin\theta_{H},\qquad$ $x_{2}=r\cos\theta_{H},\qquad$ $x_{3}=r\sin\varphi_{H}\sin\theta_{H},\qquad$ $\text{for }x\in\Omega_{H}.$
@@
$\hat{r}\vcentcolon=\frac{x}{|x|},\qquad\hat{\theta}_{V}\vcentcolon=\frac{x_{1}x_{3}\hat{e}_{1}+x_{2}x_{3}\hat{e}_{2}-(x_{1}^{2}+x_{2}^{2})\hat{e}_{3}}{|x|\sqrt{x_{1}^{2}+x_{2}^{2}}},\qquad\hat{\varphi}_{V}\vcentcolon=\frac{-x_{2}\hat{e}_{1}+x_{1}\hat{e}_{2}}{\sqrt{x_{1}^{2}+x_{2}^{2}}}.$
For $x\in\Omega_{H}$, unit vectors $(\hat{r},\hat{\theta}_{H},\hat{\varphi}_{H})$ is defined by exchanging the terms $x_{2}\leftrightarrow x_{3}$ and $\hat{e}_{2}\leftrightarrow\hat{e}_{3}$ on the above expression. Let $\chi_{V}$, $\chi_{H}\in\mathcal{C}^{\infty}(\mathbb{R}^{3})$ be non-negative functions satisfying
$0\leq\chi_{V},\,\chi_{H}\leq 1,\qquad\chi_{V}+\chi_{V}\geq 1,\qquad\text{supp}\chi_{V}\subseteq\Omega_{V},\qquad\text{supp}\chi_{H}\subseteq\Omega_{H},$ (2.23a)
$|\nabla\chi_{H}|^{2}\leq C\chi_{H},\hskip 18.49988pt|\nabla\chi_{V}|^{2}\leq C\chi_{V},\hskip 18.49988pt|D^{k}\chi_{H}|+|D^{k}\chi_{V}|\leq C,$ (2.23b)
$\hat{r}\cdot\nabla\chi_{H}=\hat{\varphi}_{H}\cdot\nabla\chi_{H}=0,\hskip 18.49988pt\hat{r}\cdot\nabla\chi_{V}=\hat{\varphi}_{V}\cdot\nabla\chi_{V}=0$ (2.23c)
for $x\in\mathbb{R}^{3}$ and $k=1,2,3,4$, where $C>0$ is a certain generic constant. The detailed construction for $\chi_{V}$ and $\chi_{H}$ is explained in Appendix A.2. For $x\in\Omega_{V}$ or $\Omega_{H}$, we define the radial and angular derivatives as follows
$\partial_{r}\vcentcolon=\frac{x}{|x|}\cdot\nabla,\hskip 18.49988pt\partial_{\theta_{V}}\vcentcolon=|x|\hat{\theta}_{V}\cdot\nabla,\hskip 18.49988pt\partial_{\varphi_{V}}\vcentcolon=\sqrt{x_{1}^{2}+x_{2}^{2}}\,\,\hat{\varphi}_{V}\cdot\nabla,$
$\partial_{\theta_{H}}\vcentcolon=|x|\hat{\theta}_{H}\cdot\nabla,\hskip 18.49988pt\partial_{\varphi_{H}}\vcentcolon=\sqrt{x_{1}^{2}+x_{3}^{2}}\,\,\hat{\varphi}_{H}\cdot\nabla.$
For simplicity, we use the following notations
$\partial\in\{\partial_{\theta_{V}},\partial_{\varphi_{V}},\partial_{\theta_{H}},\partial_{\varphi_{H}}\},\qquad D\in\{\partial_{r},\partial_{\theta_{V}},\partial_{\varphi_{V}},\partial_{\theta_{H}},\partial_{\varphi_{H}}\}.$
In what follows, we only derive estimates for one of $\Omega_{V}$ and $\Omega_{H}$ since the other follows exactly from the same argument. To simplify, we use the notations
$\chi\vcentcolon=\chi_{V}\ \text{or}\ \chi_{H},\hskip 18.49988pt\tilde{\Omega}\vcentcolon=\Omega_{V}\ \text{or}\ \Omega_{H},$ (2.24a)
$(\theta,\varphi)\vcentcolon=(\theta_{V},\varphi_{V})\ \text{or}\ (\theta_{H},\varphi_{H}),\hskip 18.49988pt\partial\in\big\{\partial_{\theta},\partial_{\varphi}\big\},\hskip 18.49988ptD\in\big\{\partial_{r},\partial_{\theta},\partial_{\varphi}\big\},$ (2.24b)
$\tilde{Q}_{T}\vcentcolon=[0,T]\times\Omega_{V}\ \text{or}\ [0,T]\times\Omega_{H},\hskip 18.49988pt\tilde{\Gamma}_{T}\vcentcolon=[0,T]\times\partial\Omega_{V}\ \text{or}\ [0,T]\times\partial\Omega_{H}.$ (2.24c)
@@
$\chi\partial^{k}\psi|_{x\in\partial\Omega}=0\qquad\text{for }\ \partial\in\{\partial_{\theta},\partial_{\varphi}\}\ \text{ and }\ k=1,2,3.$ (2.25)
@@
Suppose $(\phi,\psi)\in X(T_{0})$ is a solution to the system (2.1), where $T_{0}>0$ is the time given in (2.3)–(2.4). Then there exists a constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for arbitrary $\delta\in(0,1)$ and $T\in[0,T_{0}]$, it holds that
@@
Taking the derivative $\partial^{k}$ on equation (2.1b) and multiplying both sides of the resultant equality by $\chi\partial^{k}\phi$, we get
@@
Taking the derivative $\partial^{k}$ on equation (2.1c) and multiplying both sides of the resultant equality by $\chi\partial^{k}\psi$, we get
@@
Summing the above two equations, integrating the resultant equality in $(t,x)\in Q_{T}\vcentcolon=[0,T]\times\Omega$ with integrating by parts using boundary condtion (2.25), we have that
@@
By the property of cut-off function $\chi$ in (2.23), we have for an arbitrary $\epsilon\in(0,1)$
@@
Using Proposition A.2, we also get
@@
For the term $\mathcal{I}_{3}$, we recall the definition (2.1e). Then integrating by parts, applying Lemma 1.1 and inequality (2.5), we get for $k=1,2,3$,
@@
Applying Propositions A.1 and A.2 we obtain that
@@
Substituting the estimates for $\mathcal{I}_{1}$$\mathcal{I}_{4}$ in (2.26), we have that
@@
$\leq$ $C\big\{N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\big\}+\epsilon N^{2}(T)+\frac{C}{\epsilon}\sum_{m=0}^{k-1}\iint_{\tilde{Q}_{T}}|\nabla\partial^{m}\psi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t,$ (2.27)
where $k=1,2,3$ and $\epsilon\in(0,1)$. For an arbitrary $\delta\in(0,1)$, let $\epsilon=\delta^{4}$ and $k=1$ in (2.27). Then we have from Lemma 2.4 that
@@
Next we let $\epsilon=\delta^{2}$ and $k=2$ in (2.27). Applying Lemma 2.4 and substituting (2.28) into the resultant inequality, we get
@@
Finally, let $\epsilon=\delta$ and $k=3$ in (2.27). Applying Lemma 2.4 and substituting (2.28)–(2.29) into the resultant inequality, we get
@@
Suppose the same assumption in Lemma 2.6 holds. Then there exists a constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for arbitrary $\delta\in(0,1)$ and $T\in[0,T_{0}]$, it holds that
@@
$\iint_{Q_{T}}\chi|\partial^{k}D_{t}\phi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t\leq\frac{C}{\delta^{7}}\Big\{N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\Big\}+C\delta N^{2}(T),$
where $k=0,1,2,3$.
@@
The equation (1.1a) yields $D_{t}\phi=-\rho_{+}\textnormal{{div}}\psi+f^{0}$. Integrating this equation in $(t,x)\in Q_{T}$, applying Lemmas 2.4 and inequality (2.5), we obtain the first estimate. Next, taking the derivative $\partial^{k}$ on (2.1b) yields $\partial^{k}D_{t}\phi=-\rho_{+}\textnormal{{div}}\partial^{k}\psi-\rho_{+}[\partial^{k},\textnormal{{div}}]\psi+\partial^{k}f^{0}$. Integrating this equation in $(t,x)\in Q_{T}$, applying Lemmas 2.4, 2.6, Proposition A.2, and inequality (2.5), we obtain the second estimate. $\square$